OpenStudy (anonymous):

When a car slams on its brakes and starts to skid, it leaves skid marks on the road. Investigators use these marks to determine how fast a car was going when it started to brake. The car's speed, v, is related to the skid mark length, d, by the equation v = sqrt(32 f d), where f is the coefficient of friction determined by the material the car was skidding on. Wet asphalt has a coefficient of friction of 0.5. If a car was driving 75 feet per second and slams on its brakes, how long of a skid mark will it leave? The skid marks will be feet long

OpenStudy (anonymous):

@DavidUsa

OpenStudy (anonymous):

just plug and chug. Solve \[75=\sqrt{32 \times .5 \times d}\] for d